OpenStudy (zab505):

f your starting salary were $50,000 and you received a 4% increase at the end of every year for 15 years, what would be the total amount, in dollars, you would have earned over the first 16 years that you worked?

OpenStudy (zab505):

@ganeshie8

ganeshie8 (ganeshie8):

geometric series

ganeshie8 (ganeshie8):

start by finding out below : first term, \(a\) = ? common ratio, \(r\) = ?

OpenStudy (zab505):

a=50000, r=1.04?

ganeshie8 (ganeshie8):

excellent ! so the series is : 50000 + 50000(1.04) + 50000(1.04)^2 + ... 16 terms, eh ?

ganeshie8 (ganeshie8):

use the partial sum of geometric series formula : http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithG12.gif

OpenStudy (zab505):

=1,001,179.38?

OpenStudy (zab505):

i'm lost

OpenStudy (aum):

Plug the numbers into the geometric sum formula. \(\Large a + ar + ar^2 + .... + ar^{n-1} = a\frac{1-r^n}{1-r}\) a = 50,000 r = 1.04 n = 16 Sum = \(\Large 50,000\frac{1-1.04^{16}}{1-1.04}\) = ?

OpenStudy (zab505):

$1,091,225.56?

OpenStudy (zab505):

@aum

OpenStudy (aum):

They want only the dollar portion and it comes to $1,091,227. Yours if off by a $1 because it depends on how many decimal places you did the numerator and the denominator in the above formula.

OpenStudy (zab505):

It also asks to round your answer to the nearest whole dollar, and express your answer without using commas? So what would be the amount they want?

OpenStudy (aum):

1091227

OpenStudy (zab505):

Thank you so much!

OpenStudy (aum):

You are welcome.